High Country, Inc., produces and sells many recreational products. The company has just opened a new plant to produce a folding camp cot that will be marketed throughout the United States. The following cost and revenue data relate to May, the first month of the plant’s operation: Beginning inventory 0 Units produced 46,000 Units sold 41,000 Selling price per unit $ 76 Selling and administrative expenses: Variable per unit $ 4 Fixed (per month) $ 560,000 Manufacturing costs: Direct materials cost per unit $ 17 Direct labor cost per unit $ 10 Variable manufacturing overhead cost per unit $ 1 Fixed manufacturing overhead cost (per month) $ 782,000 Management is anxious to assess the profitability of the new camp cot during the month of May. Required:

1. Assume that the company uses absorption costing.

a. Determine the unit product cost.

b. Prepare an income statement for May.

2. Assume that the company uses variable costing.

a. Determine the unit product cost.

b. Prepare a contribution format income statement for May.
Answer #1
1a
Direct materials cost 17
Direct labor cost 10
Variable manufacturing overhead 1
Fixed manufacturing overhead 17 =782000/46000
Unit product cost 45
1b
Absorption costing Income statement
Sales 3116000 =41000*76
Cost of goods sold 1845000 =41000*45
Gross profit 1271000
Selling and administrative expenses 724000 =560000+(41000*4)
Net operating income 547000
2a
Direct materials cost 17
Direct labor cost 10
Variable manufacturing overhead 1
Unit product cost 28
2b
Contribution format income statement
Sales 3116000 =41000*76
Variable expenses:
Variable cost of goods sold 1148000 =41000*28
Variable selling and administrative expenses 164000 =41000*4
1312000
Contribution margin 1804000
Fixed expenses:
Fixed manufacturing overhead 782000
Fixed selling and administrative expenses 560000
1342000
Net operating income 462000